Sol Summary – December 27th

Crew 201 Sol Summary Report 27-Dec-2018

Sol: 12

Summary Title: This is about to come to its end.

Author’s name: Carlos Mariscal, XO.

Mission Status: Experiments are having some slow progress. Unfortunately, out VSAT, which is necessary to do one of the experiments, will not make it to the MDRS on time, so that experiment will not be performed. Depending on the progress of another of the experiments, we could end simulation tomorrow during the day or evening.

Sol Activity Summary: We:

– Continued activities of the artificial vision experiment.

– Printed several pieces for the rover.

– Had a lot of work at the GreenHab; we planted some seeds and organized some stuff.

– Recorded videos for media.

– Had a great time with some pictures in the snow (before it got kind of bad).

Look Ahead Plan: We hope to have our last EVA tomorrow to test mechanical components of the rover. Besides that, we will finish some experiments and start with the cleaning of the station for the next crew.

Anomalies in work: The plumb and water tank deicing system didn’t work properly, so we could not use any water during the morning. Fortunately, the temperature raised enough to get water up just before the snow started.

Weather: Cold in the morning, then snow, snow, wind and snow.

Crew Physical Status: A crew member had headaches, nothing serious, an aspirin solved it.

EVA: No EVA today.

Reports to be file:

– Commander report

– Sol summary

– HSO report

– Operations report

– GreenHab report

– Astronomer report

– Journalist report

– EVA request

– Photos

Support Requested: No support requested.